Personal Investment Analysis Find of the cost of a bachelor's degree at the university of your...

  1. Personal Investment Analysis Find of the cost of a bachelor's degree at the university of your choice assume additional costs of $16,000 for an additional fifth year of education to get Master's degree. Assume that all tuition is paid at the beginning of the year. A student considering this investment must evaluate the present value of cash flows from possessing a graduate degree versus holding only the undergraduate degree. Assume that the average student with an undergraduate degree is expected to earn an annual salary of $55,000 per year (assumed to be paid at the end of the year) for 10 years. Assume that the average student with a graduate Masters degree is expected to earn an annual salary of $76,000 per year (assumed to be paid at the end of the year) for nine years after graduation. Assume a minimum rate of return of 10%. Determine the net present value of cash flows from an undergraduate degree. Use the present value table provided in this chapter 26. Determine the net present value of cash flows from a Masters degree, assuming that no salary is earned during the graduate year of schooling. What is the net advantage or disadvantage of pursuing a graduate degree under these assumption? Two sources need to be cited.
Answer #1

ANSWER

NPV of undergraduate level = -X + $55,000 * (Present value annuity factor for 10 years at 10%)

= -X + $55,000 * 6.1447

= -X + 337,958.5

NPV of Graduate level;

= -x -16,000 + 66,000 (Present value annuity factor for 10 years at 10%)
= -X -16,000 + 66,000* 6.1447

= -X - 12,000 + 405,550.2

= -X + 393,550.2

NPV of graduate level is greater than the undergraduate level and thus, it is preferred to pursue graduate level.

===========================================

DEAR STUDENT,

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive ratings

*****************THANK YOU****************